System of Equation:
Andrea sells flower pot of different sizes. The price of a small flower pot is RM10, medium flower pot RM15 and large flower pot RM 40. Every months, the number of small flower pots sold is equal to the number of the medium and large flower pots sold. The number of medium flower pots sold is twice the number of the large flower pots sold. Andrea needs to pay a rent of RM300 per month for her business premise. What are the minimum numbers of pots of each size which Andrea has to sell in a month so that she can pay the monthly rent?

Answer by ikleyn(52798)

.
Andrea sells flower pot of different sizes. The price of a small flower pot is RM10, medium flower pot RM15 and large flower pot RM 40.
Every months, the number of small flower pots sold is equal to the number of the medium and large flower pots sold.
The number of medium flower pots sold is twice the number of the large flower pots sold.
Andrea needs to pay a rent of RM300 per month for her business premise. What are the minimum numbers of pots of each size
which Andrea has to sell in a month so that she can pay the monthly rent?
~~~~~~~~~~~~~


Let x be the number of the LARGE flower pots.

Then the number of the medium pots is 2x, according to the condition, and

     the number of the small spots is (2x + x) = 3x,  again, in accordance with the condition.


The total money equation is THIS

    10*(3x) + 15*(2x) + 40x = 300   monetary units.


Simplify and find x


    30x + 30x + 40x = 300,

         100x       = 300

            x       = 300/100 = 3.


ANSWER.  3 large pots;  2*3 = 6 medium pots  and  3*3 = 9 small pots.

Solved.

The remarkable fact is that the provlem is solved using only one unknown and one equation.